Find four consecutive even integers whose sum is 340
Answer by checkley77(12844)

LET X BE THE FIRST INTEGER.
X+2, X+4, X+6 BE THE SUBSEQUENT INTEGERS.
X+(X+2)+(X+4)+(X+6)=340
4X+12=340
4X=340-12
4X=328
X=328/4
X=82 FOR TRHE SMALLEST INTEGER.
82+2=84 FOR THE NEXT INTEGER.
82+4=86 FOR THE NEXT INTEGER.
82+6=88 FOR THE LARGEST INTEGER.
PROOF:
82+84+86+88=340
340=340